Please help me solve this equation, -8x +1 = 81 + 2x. Thank you.
Answer by ewatrrr(24785)


Hi
Solving for x (isolating x on one side of the equation)
-8x +1 = 81 + 2x
-80 = 10x
-8 = x
CHECKING our Answer***
-8(-8)+1 = 65 = 81- 16